OpenStudy (anonymous):

MEDAL AND FAN!!!!!!! PLEASE HELP! Divide and select the answer in simplest form: 3/8+2 1/4 A. 1/6 B. 1/5 C. 1/4 D. 1/3

OpenStudy (anonymous):

2 and 1/4

OpenStudy (anonymous):

@JimmiFizzle @KeganMoore @Aripotta

OpenStudy (anonymous):

2 and 1/4 as an improper fraction is 9/4 so its 9/4 divided by 3/8

OpenStudy (anonymous):

get a common denominator so its 18/8 and 3/8

OpenStudy (anonymous):

okay so far following

OpenStudy (anonymous):

the answer is A. 1/6

OpenStudy (anonymous):

tyhamnk you !!!1

OpenStudy (anonymous):

no wait thats wrong

OpenStudy (anonymous):

C. 1/4

OpenStudy (anonymous):

aha okay thanks

OpenStudy (anonymous):

How is it 1/4?

OpenStudy (anonymous):

oooooomyyyygoshhhhhh

OpenStudy (anonymous):

did you get my message?

OpenStudy (anonymous):

@xokatexo How did you get 1/4. It doesn't make sense to me.

OpenStudy (anonymous):

It's 1/6

OpenStudy (anonymous):

\[\frac{ 9 }{ 4 }\div \frac{ 3 }{ 8 }\] Flip the first fraction and multiply, \[\frac{ 4 }{ 9 }\times \frac{ 3 }{ 8 }\] \[\frac{ 12 }{ 72 }\] Simplify \[\frac{ 1 }{ 6 }\] Or. Change it to decimals. .375/2.25 =.16666666666667 Change it to a fraction you get 1/6.

OpenStudy (anonymous):

omg yeah sorry i did the math wrong yeah its 1/6 sorry!